OpenStudy (anonymous):

Factor completely: 4x2 + 25x + 6 (4x + 1)(x + 6) (4x + 6)(x + 1) (2x + 3)(2x + 2) (2x + 6)(2x + 1)

OpenStudy (ash2326):

@Solmyr Did you try this?

OpenStudy (anonymous):

sadly no cause i have no idea of how to do it

OpenStudy (anonymous):

i was looking for someone to show me how to do it

OpenStudy (ash2326):

Ok, I'll explain you

OpenStudy (ash2326):

General quadratic equation \[ax^2+bx+c=0\] Here we have \[4x^2+25x+6\] So what's a , b and c here?

OpenStudy (anonymous):

a=4 b=25 c=6

OpenStudy (ash2326):

Good, first step is to find the product of a and c What is it here?

OpenStudy (anonymous):

4*6=24

OpenStudy (ash2326):

Now b is 25 here Can you find factors of 24, which add up or subtract to form 24 ?

OpenStudy (anonymous):

2,4,6,8

OpenStudy (anonymous):

12

OpenStudy (ash2326):

Let me rephrase it, 2 factors of 24 which add up or subtract to form 25?

OpenStudy (anonymous):

im not sure

OpenStudy (ash2326):

Check if 24 and 1 satisfy that??

OpenStudy (anonymous):

ok i guess it does

OpenStudy (ash2326):

so write 25x as 24x and 1x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have to do something ill brb

OpenStudy (ash2326):

\[4x^2+24x+x+6=0\]

OpenStudy (anonymous):

im back